OpenStudy (anonymous):

Gino designed a garden for his rectangular backyard. The design he drew on paper is 10 inches long. The length of his backyard is 40 feet. If his drawing is to scale, and his backyard is 28 feet wide, how many inches wide is his drawing? Write a proportion using words. Then, use the proportion to calculate the drawing width. Show your work.

OpenStudy (anonymous):

since the drawing can be scaled and the ratio of a feature from drawing:garden is coonstant, we can write, constant scale ratio=10/40=1/4=drawing length/actual length similarly writing drawing width/actual width=1/4 and so we get drawing width=7